OpenStudy (anonymous):

WILL BE GIVING MEDALS FOR RIGHT ANSWER! PLEASE! Which function below is the inverse of f(x) = x^2 - 6? A: x^2/16 B: +-4sqrtx C:+-sqrtx+16 1/x^2-16

OpenStudy (mathstudent55):

Do you know how to find the inverse of a function?

OpenStudy (anonymous):

Not one with a exponent. @mathstudent55

OpenStudy (anonymous):

@thomaster

OpenStudy (mathstudent55):

The important thing is to know the method. Here is the method, then we'll go over the problem step by step. 1. Write y instead of f(x) 2. Switch x and y. 3. Solve for y. 4. Replace y with \(f^{-1}(x) \)

OpenStudy (anonymous):

Well, if you have f(x)= x^2 - 6 you would switch them to y=x^2 -6 x=y^2 -6 and solve for y. so x+6= y^2 To cancel the square you would squareroot each side. so the \[\sqrt{x+6} = y\]

OpenStudy (anonymous):

So the problem would look like this X = y^2 - 6? Thank you @AJBB I need to know how to do this problem also though. Like really thoroughly.

OpenStudy (anonymous):

That isn't even one of the answers either. @AJBB

OpenStudy (mathstudent55):

\( f(x) = x^2 - 6 \) 1. \( y = x^2 - 6 \) 2. \( x = y^2 - 6 \) 3. \( x = y^2 - 6 \) \(x + 6 = y^2\) \(y^2 = x + 6\) \(y = \pm \sqrt{x + 6} \) 4. \(f^{-1}(x) = \pm \sqrt{x + 6} \)